Two resistors connected in series have an equivalent resistance of 699.9 Ω. When they are connected in parallel, their equivalent resistance is 124.3 Ω. Find the resistance of each resistor.

Ω (small resistance)

Ω (large resistance)
Answer #1

Surius Pasalle R,R124.3 R, г - 124.3 K,R2 699.9x1243 R/R2; 86997:57 Ri-R2 316.662 R 538 28n R2 - 699.7 S332161.622-
Answer #2

SOLUTION :


2 resistors, R1 and R2 connected in series = 699.9 ohms

=> R1 + R2 = 699.9 ohms

=> R2 = 699.9 - R1


When they are connected in parallel, the equivalent is = 124.3 ohms 

=> 1/ [1/R1 + 1/R2] = 124.3 ohms

=> R1 R2 / [R1 + R2] = 124. 3 ohms

Substituting value of R2 = 699.9 - R1 and R1 + R2 = 699.9 :

=> R1(699.9 - R1) / 699.9 = 124.3 

=> 699.9 R1 - R1^2 = 124.3 * 699.9

=> R1^2 - 699.9 R1 + 86997.57 = 0 

=> R1 = 699.9/2 +/- sqrt(699.9^2 - 4(1)(86997.57)) / 2

=> R1 = 349.95 +/- 188.33

=> R1 = 161.62, 538 28 

So, if R1 = 161.62, R2 = 699.9 - 161.62 = 538.28 and vice-versa.


Hence, 


R1 = 161.62 ohms and R2 = 538.28 ohms 

OR

R1 = 538.28 ohms and R2 = 161.62 ohms.


It means one restore is 161.62 ohms and the other resistor is 538.28 ohms 

(ANSWER)


answered by: Tulsiram Garg